### v3

# Spatiotemporal Photon Blockade for Nonreciprocal Quantum Absorption

Abstract
Controlling the flow of photons is crucial for advancing quantum technologies. We introduce the concept of spatiotemporal photon blockade for nonreciprocal quantum absorption, utilizing space-time-periodic metasurfaces. Our study presents a methodology for experimentally realizing this effect, where photon frequency coherence with the metasurface’s space-time modulation enables one-way quantum absorption. In this system, forward-traveling photons are energetically modulated and absorbed within the slab, while backward-traveling photons are transmitted without interaction. Our analysis includes band structure, isofrequency diagrams, and nonreciprocal absorption results. These findings lay the groundwork for developing nonreciprocal quantum devices and enhancing photon management in milli-Kelvin temperature quantum systems.
Photon blockade is a quantum optical phenomenon analogous to the Coulomb blockade, where the suppression of electrical transport occurs due to electronic repulsion, preventing multiple electrons from occupying the same region. In photon blockade, the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons. This phenomenon occurs in systems where strong nonlinear interactions between photons effectively create a scenario in which only one photon can be absorbed or transmitted at a time, leading to quantized light behavior. This effect, typically observed in systems such as atoms, quantum dots, or superconducting circuits coupled to resonant cavities, involves the system entering an excited state after the first photon interaction. This state shift prevents further photon interactions until the system returns to its ground state, which is crucial for quantum information processing and communication. Historically, photon blockade has been achieved in systems with strong coupling between quantum emitters and optical cavities, where the nonlinearity creates an energy gap sufficient to block additional photons  (?, ?). Recent advancements have expanded this concept to nonreciprocal photon blockade in rotating nonlinear devices, where device geometry or external fields control the directionality of photon interactions  (?). Additionally, loss-induced nonreciprocal photon blockade has been demonstrated, where engineered losses enable nonreciprocal photon dynamics, enhancing control over photon flow in quantum circuits  (?, ?).
Quantum nonreciprocity is increasingly relevant for superconducting circuits, which are essential for scalable quantum computing and communication networks. Nonreciprocal elements like isolators and circulators are critical for protecting qubits from back-reflected signals and routing quantum information. Recent studies have explored various mechanisms for achieving quantum nonreciprocity, including dynamic modulation and parametric interactions in superconducting devices  (?, ?). Observations of superconducting nonreciprocity through diamagnetic effects and engineered circuit asymmetries offer robust platforms for nonreciprocal quantum devices  (?, ?, ?, ?). Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of quantum technologies. To overcome these challenges, we propose a novel approach for nonreciprocal photon blockade through spatiotemporal modulation in quantum metasurfaces. By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional quantum absorption, where photons are selectively absorbed or transmitted based on their direction of incidence. This spatiotemporal photon blockade presents new opportunities for controlling quantum light-matter interactions and offers significant advancements for quantum information technologies and nonreciprocal photonic devices. Our approach achieves efficient directional absorption with simultaneous amplification, a rare combination in conventional devices. It addresses inefficiencies and spurious signal generation, paving the way for high-performance signal processing applications.
Space-time refractive index modulation is conventionally realized through the modulation of electrical permittivity, as it is challenging to spatio-temporally modulate the magnetic permeability of materials  (?). In  (?, ?), it was shown for the first time that modulating both permeability and permittivity results in physical phenomena that cannot be achieved through conventional permittivity-modulated media, including zero reflection, zero band-gap, a large nonreciprocal response, and strong space-time harmonics. Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and time-varying systems (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). Utilizing analytical techniques such as Bloch-Floquet solutions, we provide a comprehensive understanding of the wave scattering phenomena from our metasurface, validating its design and offering insights for further optimization. Unlike linear systems constrained by phase-matching and dispersion, our nonlinear spatiotemporal media achieve superior performance in a compact form, making them ideal for quantum technologies.
1 Spatiotemporal Photon Blockade
Consider photons with frequency $\omega_{0}$ incident on a slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$. We demonstrate that the temporal coherence between the incident photons and the periodic modulation of the superconducting slab induces a nonreciprocal photon blockade. This phenomenon arises because the energy associated with the space-time modulation effectively inhibits the transmission of subsequent photons. Instead, the incident photons transition to higher energy states within the slab and are absorbed, preventing their exit. This results in a one-way quantum absorption process, facilitated by the coherent spatiotemporal modulation of the superconducting slab.
Figure 1: Nonreciprocal photon blockade in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident photons share the same frequency ($\omega_{\text{s}}=\omega_{0}$). The system comprises space-time-modulated nonlinear unit cells coupled together. a) Incidence from the left port ($L_{\text{in}}$): Photon blockade and energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$). b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of photons to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of photons.
Figure 1 illustrates the principle of spatiotemporal photon blockade enabling nonreciprocal quantum absorption by leveraging the interplay between the incident photons and the one-way coherent dynamic modulation. Figure 1a demonstrates the photon blockade effect when photons are incident from the left port ($L_{\text{in}}$). Here, the photons interact strongly with the space-time modulation traveling from left to right. This interaction causes the system to enter an excited state, leading to energy transitions and blocking the transmission of photons to the right port ($R_{\text{out}}=0$). In contrast, Fig. 1b illustrates the scenario for photons incident from the right port ($R_{\text{in}}$). In this case, the interaction with the space-time modulation traveling in the opposite direction is weak. Consequently, there is minimal energy transition and the photons are transmitted freely to the left port ($L_{\text{out}}=R_{\text{in}}$). This nonreciprocal behavior, where the photon transmission and absorption are direction-dependent, underscores the potential of spatiotemporal metasurfaces for applications in quantum communication and sensing.
Figure 2: One-way photon blockade and absorption functionality.
Figure 2 illustrates the proposed one-way photon blockade and absorption mechanism facilitated by the coherently driven quantum spatiotemporal metasurface and its functionality. The metasurface operates as a nonreciprocal element, permitting photon transmission in one direction while blocking it in the opposite direction. To understand the nonreciprocal behavior of the spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation. The analysis reveals crucial insights into the energy transitions and nonreciprocity induced by the space-time modulation. Figure 3a presents the $\omega-\kappa$ diagram, showcasing the band structure of the metasurface at the critical point where $\omega_{0}/\omega_{\text{s}}=1$. This diagram illustrates a strong nonreciprocity in the system. Specifically, the higher-order harmonics that travel in the forward $+z$ direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$. This convergence indicates a resonant interaction between the incident photons and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic. The absorbed energy causes a transition from the fundamental frequency $\omega_{0}$ to higher energy states denoted as $\omega_{m}$.
Figure 3: Band structure of the spatiotemporal quantum metasurface computed. (a) The $\omega-\kappa$ diagram illustrates strong nonreciprocity at $\omega_{0}/\omega_{\text{s}}=1$. Here, all higher-order forward ($+z$ direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$. In contrast, the higher-order backward ($-z$ direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states. (b) The $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$ further demonstrates the nonreciprocity of the metasurface, where all forward higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$, effectively absorbing the energy of the fundamental harmonic. The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the $+z$ direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
In contrast, the higher-order harmonics that travel in the backward $-z$ direction are spatially separated in the $\omega-\kappa$ diagram. This separation prevents any significant interaction between the incident photons and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states. As a result, the backward-propagating waves do not experience the same degree of energy absorption as the forward-propagating waves, which directly contributes to the observed nonreciprocal behavior of the metasurface. Figure 3b shows the $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$. This diagram further confirms the nonreciprocal nature of the metasurface, where all forward-propagating higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$. The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the $+z$ axis, parallel to the metasurface boundary. This confinement prevents the transmission of energy to the exterior of the metasurface, leading to strong nonreciprocal absorption.
The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade. In photon blockade, the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission. In our system, this blockade effect is realized through the spatiotemporal modulation of the metasurface. The space-time modulation, with a frequency $\omega_{\text{s}}$ matched to the incident photon frequency $\omega_{0}$, induces a strong nonlinear interaction for forward-propagating waves. As a result, when photons from the left are incident on the metasurface, they are resonantly coupled to the modulation, leading to energy absorption and transition to higher frequency states (spatial harmonics) $\omega_{m}$. This absorption effectively prevents the transmission of these photons through the metasurface, analogous to a photon blockade. The nonreciprocity arises because the same modulation does not facilitate similar energy transitions for backward-propagating waves, allowing them to pass through the metasurface without significant energy loss. Thus, the metasurface behaves as a nonreciprocal photon blockade device, where the space-time modulation selectively blocks photons traveling in one direction by shifting their energy to higher states, while allowing photons traveling in the opposite direction to pass freely. This behavior is crucial for potential applications in quantum information processing and nonreciprocal quantum devices, where controlling the directionality and energy states of photons is essential.
Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal quantum metasurface. The metasurface is designed to exhibit strong nonreciprocal behavior due to the space-time modulation, leading to direction-dependent absorption and transmission properties. Figures 4a and 4b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles. Specifically, Figure 4a demonstrates the field distribution and frequency response when an electromagnetic wave is incident from the left at an angle of $55^{\circ}$. The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission, indicating effective photon blockade in this direction. The corresponding frequency spectrum confirms the absorption by showing the lack of transmitted energy at the incident frequency. Conversely, Figure 4b presents the scenario where the incident beam approaches from the right at an angle of $125^{\circ}$. In this case, the metasurface allows full transmission of the wave to the left side, effectively reflecting the wave back towards the source at an angle of $55^{\circ}$. The frequency spectrum corroborates this by displaying a significant transmitted signal at the original frequency, confirming the nonreciprocal behavior of the metasurface. Figures 4c and 4d provide the magnetic field distribution ($H_{z}$) under the same conditions. Figure 4c shows the absorption of the left-incident wave at $55^{\circ}$, while Figure 4d demonstrates the full transmission of the right-incident wave at $125^{\circ}$. The consistent behavior observed in both the electric and magnetic field distributions confirms the robustness of the nonreciprocal absorption mechanism induced by the spatiotemporal modulation in the superconducting quantum metasurface. These results demonstrate the capability of the designed metasurface to enforce nonreciprocal photon blockade, effectively controlling the directionality of absorption and transmission. The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of photon propagation.
Figure 4: Nonreciprocal absorption in spatiotemporal quantum metasurfaces. (a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$. (c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.